# Synthetic fractional flux quanta in a ring of superconducting qubits

### v1

@@ Abstract
In Fig. 3a) we present the results for weak driving $\Omega=0.01~{}{\rm GHz}$. A clear broad bright signal following the dispersion $E^{(1)}_{k}=\omega-\omega_{d}-J\cos(2\pi k/N_{s}+\Phi/N_{s})$, for $k=0,\ldots,N_{s}-1$ emerges. This feature, displaying the full $2\pi$ flux periodicity, is identified as corresponding to the $N_{p}=1$ sector. Remarkably, a weaker signal in the form of a dip is visible, in lower part of Fig. 3a), that corresponds to the $N_{p}=2$ sectors. Here, we see the inception of the reduced periodicity, where a new minimum appears between $\Phi=0$ and $\Phi=2\pi$. The apparent deviations with respect to the perfect reduced periodicity were demonstrated to arise as $1/N_{s}$ corrections [63, 10, 13]. In order to observe the $N_{p}=3$ bound state we needed to increase the driving amplitude $\Omega$ to be of order $J$. We comment that the time scale for photon hopping from the transmission line to the lattice site $j=1$, determined by $2\pi/\Omega$, has to be comparable to or shorter than the time scale on which photons hop from one transmon to the other, $1/J$, requiring $\Omega\gtrsim J$. For $\Omega=0.04~{}{\rm GHz}$ the sector $N_{p}=3$ appears, as shown in Fig. 3b). The lowest energy band is displayed as significantly broadened and we can only observe the fractional $32\pi/$ 3 periodicity in the excited bands. We also point out that degeneracies in the level crossings of the many-body spectrum, taking place at specific values of the synthetic flux, are lifted as a combined effect of driving and dissipation.
@@
with $E_{0}^{(N_{p})}=N_{p}(\omega-\omega_{d})-\frac{U}{2}N_{p}(N_{p}-1)$ and $J_{(N_{p})}=\frac{JN_{p}}{(N_{p}-1)!}(J/U)^{N_{p}-1}$ [64]. This formula is in good agreement with the exact spectrum shown in Figs. 3c) and d).
@@
In summary, our analysis shows that the low energy bands in the absorption spectrum have a periodicity reduced by $N_{p}$ reflecting the size of the bound state formed in the system. According to the predictions in Ref.[10], for sufficiently strong attractive interactions the energy bands corresponding to bound states detach from the single particle (scattering) band [see Fig. 3c)]. We note that solitons are increasingly more fragile with $N_{p}$ to relaxation and dephasing. This is due to the fact that the solitons take approximately a time $\sim N_{s}/J_{(N_{p})}$ to go around the ring of transmons. Therefore, to observe coherent flux oscillations such a time has to be shorter than the single transmon relaxation time $\sim 1/N_{p}\gamma_{1}$ and pure dephasing time $1/\gamma_{\phi}$ [65]. Thus we require $N_{p}\gamma_{1},\,\gamma_{\phi}\ll J_{(N_{p})}/N_{s}$ (see [39] for an assessment of the coherence properties).
@@
Conclusions.— We have shown how a ring-shaped system of capacitively-coupled transmons can be employed to study the many-body spectrum of a strongly correlated bosonic system in presence of an effective magnetic field. The latter is obtained through a suitable Floquet driving protocol employing Lorentzian pulses (Levitons). Such a scheme provides a new interesting scenario in the field of driven quantum system. This way we widen the scope of the superconducting circuits quantum simulators significantly.
The present implementation employing transmons can access the physics of the system in regimes that are difficult to attain with other platforms, for instance cold atoms. In particular, we focus here on the phenomenon of fractionalization of flux quantum that is predicted to be displayed in the ‘band structure’ of the system (energy versus magnetic field) due to the formations of lattice bright solitons [10, 13]: because of the formation of the bound state of effective mass $N_{p}$, the flux quantum is reduced by the same factor $N_{p}$. Therefore, the effective periodicity of the energy versus applied flux is also reduced and the persistent current displays a fractional periodicity. With our work, we have demonstrated, for realistic experimental conditions  [17], that fractional flux quanta are indeed visible via the reflection protocol we propose, see Fig. 3. Important problems as the impact of disorder on the fractionalization, macroscopic phase coherence and SQUID physics [30], Aharonov-Bohm oscillations in bosonic systems [66], $N$-component bosons, and mesoscopic simulation of lattice gauge theories [67], can be studied in a regime of parameters that is complementary to the working conditions of cold atoms systems.
Acknowledgements.— We thank B. Blain, E. C. Domanti, and G. Marchegiani for discussions.
@@ I The system Hamiltonian
The Lagrangian ${\cal L}$ of the circuit is a function of the $\phi_{j}$ and their time derivative $\dot{\phi}_{j}$, (${\cal L}(\{\phi_{j}\},\{\dot{\phi}_{j}\})$, and it is given by the difference between the charging energy $K$ and the Josephson potential $U$,
@@
${\cal C}^{-1}_{i,j}=\frac{z_{-}^{|i-j|}}{C_{0}(z_{+}-z_{-})}=\frac{e^{-|i-j|/\ell_{s}}}{2C_{0}\sqrt{2\eta+\eta^{2}}}$ (S6)
@@
$H_{j}=4E_{C,j}N^{2}_{j}-E_{J,j}\cos(\phi_{j})$ (S7)
where $N_{j}=q_{j}/2e=-i\partial/\partial\phi_{j}$ is the number operator conjugate to the phase, $[\phi_{j},N_{j}]=i$. The transmons have in principle all slightly different Josephson energy $E_{J,j}$ and charging energy $E_{C,j}=e^{2}/(2(C_{J,j}+C_{g,j}))$. Nevertheless, each Josephson junction can be realized as a parallel of two nominally equal junctions (SQUID), that can be tuned through external fluxes both in a static and time-dependent way. We then assume the transmon to be nominally equal and introduce bosonic operators describing plasmonic modes $\phi_{j}=\frac{\ell}{\sqrt{2}}(a_{j}+a^{\dagger}_{j})$ and $N_{j}=\frac{-i}{\ell\sqrt{2}}(a_{j}-a^{\dagger}_{j})$, with $\ell=(8E_{C}/E_{J})^{1/4}$. By expanding the Josephson potential up to fourth order in the phases and retaining only number-conserving terms, we obtain the Hamiltonian
@@
I.1 Time-dependent modulation
@@ II Leviton pulses
${\cal I}_{1}$ $=$ $1-\frac{2\sinh(2\pi\tau/T)\sin(\pi\delta t/T)}{\sin(\pi(\delta t+2i\tau)/T)},$ (S34)
@@
${\cal I}_{1}\simeq e^{2\pi i\delta t/T}\left[1-4e^{-4\pi\tau/T}\sin^{2}(\pi\delta t/T)\right].$ (S36)
@@
II.1 Coupling to Floquet side bands
@@
$J^{\prime}_{(n)}/J_{0}^{\prime}$ $=$ $\int_{0}^{T}\frac{dt}{T}e^{2\pi int/T}e^{if(t)}$ (S39)
$=$ $\int\frac{dz}{2\pi i}z^{n-2}\frac{1-\kappa z}{1-\kappa/z}$ (S40)
$=$ $\begin{cases}\kappa^{n-1}-\kappa^{n+1},&\text{if $n\geq 1$},\\ -\kappa,&\text{if $n=0$},\\ 0&\text{if $n<0$}\end{cases}$ (S41)
@@
$J_{(n)}/J_{0}$ $=$ $\int_{0}^{T}\frac{dt}{T}e^{2\pi int/T}e^{if(t)-if(t-\delta t)}$ (S42)
$=$ $\frac{1}{w}\int_{C}\frac{dz}{2\pi i}z^{n-2}\frac{(1-\kappa z)(z-\kappa w)}{(1-\kappa/z)(1-\kappa z/w)}$
$=$ $\frac{1}{w}\sum_{p,q=0}^{\infty}\frac{\kappa^{p+q}}{w^{q}}\left[-\kappa w\delta_{p,q+n-1}\right.$
$+$ $\left.(1+\kappa^{2}w)\delta_{p,q+n}-\kappa\delta_{p,q+n+1}\right],$ (S44)
with $w=e^{2\pi i\delta t/T}$. In particular, we have that
$\sum_{p,q=0}^{\infty}\frac{\kappa^{p+q}}{w^{q}}\delta_{p,q+n-r}=\begin{cases}\frac{\kappa^{n-r}}{1-\kappa^{2}/w}&\text{for $n\geq r$}\\ \frac{(\kappa/w)^{r-n}}{1-\kappa^{2}/w}&\text{for $n<r$}\end{cases}.$ (S45)
The coupling mediated by $J_{(n)}$ is always perturbatively small in $\kappa$, so that it can be safely neglected. In turn, the coupling to the first Floquet band for the rescaled case $J_{0}^{\prime}=J_{0}/\kappa$ is of order $1/\kappa$, so that the first Floquet band is accessible to the dynamics and the corrections are on order $TJ_{0}^{2}/\kappa^{2}$. It follows that the first Floquet band can be neglected for $T\ll\kappa^{2}/J_{0}$, in general a more restrictive condition that the one given in Eq. (S38).
II.2 Working at the sweet spot of flux noise insensitivity.
@@ II.3 Source signal
[…]
@@
II.4 Full time-dependent dynamics
In order to check the prediction of the Floquet theory we check the result for the effective model in fundamental Floquet band and the exact dynamics generated by the time-dependent modulation against the target evolution with constant uniform hopping and a flux $\phi$. In order to keep the analysis simple, we restrict to the single-particle spectrum and solve the dynamics for a set of $N_{s}=4$ transmon. In this subspace, the target Hamiltonian in the frame rotating at frequency $\omega$ reads
@@ II.4 Full time-dependent dynamics
where we have gauged the dependence on the external flux between the second and third site. In turn, the evolution in the fundamental Floquet band is described by the effective Hamiltonian
@@
In order to test the goodness of the Floquet modulation, we also study the fidelity $F=|\langle\psi_{\rm target}|P|\psi_{\rm exact}\rangle|$ integrated over the time/phase-shift as a function of time for different values of $T$, while keeping constant ratio $\tau/T=0.3$. The matrix $P={\rm diag}(1,-1,-1,1)$ accounts for the $\pi$ phase shifts due to the Levitons in the external links, as in Eq. (S35). As expected the larger the Floquet frequency $2\pi/T$ the better is the fidelity, as shown in Fig. (S2) where we plot the infidelity $1-F$ for four different values of $T$. For $T=1~{}{\rm ns}$ we see that the sampling at half period shows oscillations at the Floquet frequency. For higher values of the Floquet frequency we see that the error accumulated grows to few percents after 10 periods of the coherent oscillations set by $J_{0}$.
@@
III Perturbative approach to the many-body spectrum
@@
IV Master equation numerical approach
@@
Figure S4: Spectrum $|S(\omega_{d})|^{2}$ as a function of the driving frequency $\omega_{d}$ and the flux $\Phi$ for $\Omega=0.04~{}{\rm GHz}$ in the region frequency around the single-photon and two-photon resonances. The other parameters are as given in the main text and the color bar is made non-linear by setting $z(x,y)^{2/5}$. Note that the range of values covered by the color scale is much smaller than in Fig. 3 (see also Fig. S5).
@@
where $\{|n\rangle\}$ is a basis of the Hilbert space of the Hamiltonian $H$. The procedure of vectorialization then amounts to a tensor product of the original Hilbert space states with itself, that clearly squares its dimension. Given an analogous matrix representation of the non-Hermitian Hamiltonian in a basis of states $\{|n\rangle\}$, $K=\sum_{n,m}K_{n,\tilde{m}}|n\rangle|\tilde{m}\rangle$, the element $(n,\tilde{m})$ of the map can be written as
@@
For the numerical simulations we work with a Fock state representation of the boson states. The Lindbladian ${\cal{L}}$ can be cast in a matrix form by vectorializing the density matrix and the steady state is given by the eigenstate corresponding to the zero eigenvalue, within numerical precision. The dimension $L$ of the Hilbert space and hence size of the matrix representation of the Hamiltonian $L\times L$ grows quickly with the number of sites $N_{s}$ and maximum number of excitations $N_{p}$, $L=\sum_{n=0}^{N_{p}}\binom{n+N_{s}-1}{n}$. Correspondingly, the size of the Lindbladian matrix grows as $L^{2}\times L^{2}$.
@@
To observe flux fractionalization in the $N_{p}=2$ and $N_{s}=3$ photon sectors we assume low transmon relaxation rate, $\gamma_{1}=1~{}{\rm MHz}$, low coupling rate to the transmission line resonator, $\Gamma=1\div 10~{}{\rm MHz}$ and zero transmon pure dephasing rate, $\gamma_{\phi}=0$. As pointed out in the main text, we expect the oscillations in the $N_{p}$ sector soliton to be visible as long as $N_{p}\gamma_{1},\gamma_{\phi}\ll J_{(N_{p})}/N_{s}$. In Fig. S3 we show the fading of the oscillations as we increase the relaxation and dephasing rates. The effective hopping for the $N_{p}=2$ soliton with $J=0.041~{}{\rm GHz}$ and $U=0.188~{}{\rm GHz}$ is $J_{(2)}=0.018~{}{\rm GHz}$, so that in the left panels, where $J_{(2)}/3>2\gamma_{1},\gamma_{\phi}$ oscillations are visible, whereas in the right panels, where $J_{(2)}/3<2\gamma_{1},\gamma_{\phi}$ oscillations are lost.
@@
In the main text we show the results of the simulations for different driving strength $\Omega$. In Fig. S4, we complement that description by showing the plot of the signal for driving $\Omega=0.04~{}{\rm GHz}$ in the region around the single-photon peak. We see a signal that is strongly modified with respect to the weak driving one. In Fig. S5, top panel, we plot the output signal $S(\omega_{d})$ as a function of the driving frequency $\omega_{d}$ and the driving strength $\Omega$ at zero flux, with specific cuts at different $\Omega$ in the bottom panel. We clearly see that the signal for increased driving strength $\Omega$ is highly modified from the weak driving case. In particular, the single-photon peak splits and transfers its spectral weight to higher photon resonances, that are located at the expected many-body energies. We then see that the two-photon and three-photon peaks are very narrow resonances for weak relaxation rate $\Gamma$, surrounded by larger dips. The peaks are washed out by increasing $\Gamma$, and only dips remain, that nevertheless show the expected evolution with the unperturbed many-body energies, as shown in the main text. In the next section we study the single site driven-dissipative dynamics in presence of Hubbard interaction and ascribe to the combination of driving, dissipation, and interaction the complicate spectrum evolution.
V Single site driven-dissipative dynamics
@@
Figure S5: Top: reflection spectrum as a function of the driving frequency $\omega_{d}$ and the driving amplitude $\Omega$, for $\omega=3.9$ GHz, and the other parameters as in Fig. 3 in the main text.
@@
[…]
@@
[…]
@@
We see that in the limit for $U\to 0$ the output field asymptotically approaches the coherent state solution $\langle a^{\dagger}\rangle_{0}=-\Omega/(\omega+i\Gamma/2)$ of a driven-dissipative harmonic oscillator. In Fig. S6 we plot the output signal as function of the driving strength $\Omega$ and the relaxation rate $\Gamma$. In the top panel we see that as we increase the driving strength the main single-particle peak splits and other peaks at the two-photon and three-photon resonances appear. As the spectral weight is gradually transferred from the single-photon peak to the other resonances a complicate response featuring dips and splittings appears, that is totally due to the combination of driving and interaction. In the bottom panel we show the evolution of the output signal by increasing the relaxation rate. A general intuitive trend appears, in which fine structure details of the signal are washed out by increasing $\Gamma$.
@@
As a check, we have confirmed that the results of Eq. (V) and the results of the numerical calculation of the steady state for a single site driven-dissipative Bose-Hubbard model coincide.

### v2

@@ IV Spectrum measurement
In Fig. 3(a) we present the results for weak driving $\Omega=0.01~{\rm GHz}$, $\Gamma=5~{\rm MHz}$, $\gamma_{1}=1~{\rm MHz}$, and we neglect $\gamma_{\phi}$ (see [46] for a discussion of the role of $\gamma_{\phi}$). Notice that typical decoherence rates are dominated by relaxation processes and that the chosen value of $\gamma_{1}$ is a worst-case scenario rather than an optimistic value. A clear broad bright signal following the dispersion $E^{(1)}_{k}=\omega-\omega_{d}-J\cos(2\pi k/N_{s}+\Phi/N_{s})$, for $k=0,\ldots,N_{s}-1$ emerges. This feature, displaying the full $2\pi$ flux periodicity, is identified as corresponding to the $N_{p}=1$ sector. Remarkably, a weaker signal in the form of a dip is visible, in lower part of Fig. 3(a), which corresponds to the $N_{p}=2$ sector. Here, we see the inception of the reduced periodicity, where a minimum appears between $\Phi=0$ and $\Phi=2\pi$. The apparent deviations with respect to the perfect reduced periodicity were demonstrated to arise as $1/N_{s}$ corrections [74, 18, 21]. In order to observe the $N_{p}=3$ bound state we needed to increase the driving amplitude $\Omega$ to be of order $J$. We comment that the time scale for photon hopping from the transmission line to the lattice site $j=1$, determined by $2\pi/\Omega$, has to be comparable to or shorter than the time scale on which photons hop from one transmon to the other, $1/J$, requiring $\Omega\gtrsim J$. For $\Omega=0.04~{\rm GHz}$, $\Gamma=0.1~{\rm MHz}$, and $\gamma_{1}=10^{-2}~{\rm MHz}$ the sector $N_{p}=3$ appears, as shown in Fig. 3(b). The lowest energy band is displayed as significantly broadened and we can only observe the fractional $2\pi/3$ periodicity in the excited bands. We also point out that degeneracies in the level crossings of the many-body spectrum, taking place at specific values of the synthetic flux, are lifted as a combined effect of driving and dissipation.
@@
with $E_{0}^{(N_{p})}=N_{p}(\omega-\omega_{d})-\frac{U}{2}N_{p}(N_{p}-1)$ and $J_{(N_{p})}=\frac{JN_{p}}{(N_{p}-1)!}(J/U)^{N_{p}-1}$ [75]. This formula is in good agreement with the exact spectrum shown in Figs. 3(c) and (d).
@@
In summary, our analysis shows that the low energy bands in the absorption spectrum have a periodicity reduced by $N_{p}$ reflecting the size of the bound state formed in the system. According to the predictions in Ref. [18], for sufficiently strong attractive interactions the energy bands corresponding to bound states detach from the single particle (scattering) band [see Fig. 3c)]. We note that solitons are increasingly more fragile with $N_{p}$ to decoherence. This is due to the fact that the solitons take approximately a time $\sim N_{s}/J_{(N_{p})}$ to go around the ring of transmons. Therefore, to observe coherent flux oscillations such a time has to be shorter than the single transmon relaxation time $\sim 1/(N_{p}\gamma_{1})$ and pure dephasing time $1/\gamma_{\phi}$ [76]. Thus we require $N_{p}\gamma_{1},\,\gamma_{\phi}\ll J_{(N_{p})}/N_{s}$, so that the protocol needs a ratio $J/U$ not too small. The chosen value of the relaxation rate in Fig. 3(b) is within the state-of-the-art values [77, 78] and we have checked that the solitons and their periodicity are visible up to $\gamma_{1}=0.1~{\rm MHz}$ (see [46] for an assessment of the coherence properties).
V Conclusions
@@
We have shown how a ring-shaped system of capacitively-coupled transmons can be employed to study the many-body dynamics of attracting bosons in presence of an effective magnetic field. The latter is obtained through a suitable Floquet driving protocol employing Lorentzian pulses (Levitons). This way, we widen the scope of the superconducting circuits quantum simulators significantly.
The bosonic effective dynamics is indeed characterized by the formation of bright solitons whose smoking gun is the fractionalization of flux quantum that is predicted to be displayed in the ‘band structure’ of the system (energy versus magnetic field) [18, 21]. Such flux quantum fractionalization emerges for realistic experimental conditions  [35], in a suitable reflection protocol (see Fig. 3).
Important problems such as the impact of disorder on the fractionalization, macroscopic phase coherence and SQUID physics [37], Aharonov-Bohm oscillations in bosonic systems [79], $N$-component bosons, and mesoscopic simulation of lattice gauge theories [80], can be studied in a regime of parameters that is complementary to the working conditions of strongly interacting cold atoms systems. Such a platform is also complementary to others describing synthetic gauge fields in weakly-coupled photons, such as those occurring in topological photonics [81]. Relying on the relevance of solitons in quantum metrology [82, 83, 84, 85, 19], we expect our scheme can open new perspectives for superconducting circuits-based quantum sensing.
VI Acknowledgements
We thank B. Blain, E. C. Domanti, and G. Marchegiani for discussions.
@@ VI Acknowledgements
[…]
@@
${\cal C}^{-1}_{i,j}=\frac{z_{-}^{|i-j|}}{C_{0}(z_{+}-z_{-})}=\frac{e^{-|i-j|/\ell_{s}}}{2C_{0}\sqrt{2\eta+\eta^{2}}},$ (S6)
@@
$H_{j}=4E_{C,j}N^{2}_{j}-E_{J,j}\cos(\phi_{j}),$ (S7)
where $N_{j}=q_{j}/2e=-i\partial/\partial\phi_{j}$ is the number operator conjugate to the phase, $\phi_{j},N_{j}]=i$. The transmons have in principle all slightly different Josephson energy $E_{J,j}$ and charging energy $E_{C,j}=e^{2}/(2(C_{J,j}+C_{g,j}))$. Nevertheless, each Josephson junction can be realized as a parallel of two nominally equal junctions (SQUID), that can be tuned through external fluxes both in a static and time-dependent way. We then assume the transmon to be nominally equal and introduce bosonic operators describing plasmonic modes $\phi_{j}=\frac{\ell}{\sqrt{2}}(a_{j}+a^{\dagger}_{j})$ and $N_{j}=\frac{-i}{\ell\sqrt{2}}(a_{j}-a^{\dagger}_{j})$, with $\ell=(8E_{C}/E_{J})^{1/4}$. By expanding the Josephson potential up to fourth order in the phases and retaining only number-conserving terms, we obtain the Hamiltonian
@@
SI.1 Time-dependent modulation
@@
${\cal I}_{1}$ $=$ $1-\frac{2\sinh(2\pi\tau/T)\sin(\pi\delta t/T)}{\sin(\pi(\delta t-2i\tau)/T)},$ (S34)
@@
${\cal I}_{1}\simeq e^{-2\pi i\delta t/T}\left[1-4e^{-4\pi\tau/T}\sin^{2}(\pi\delta t/T)\right].$ (S36)
@@
SII.1 Coupling to Floquet side bands
@@
$J^{\prime}_{(n)}/J_{0}^{\prime}$ $=$ $\int_{0}^{T}\frac{dt}{T}e^{2\pi int/T}e^{-if(t)}$ (S39)
$=$ $\int\frac{dz}{2\pi i}z^{n-1}\frac{1-\kappa z}{z-\kappa}$ (S40)
$=$ $\begin{cases}\kappa^{n-1}-\kappa^{n+1},&\text{if $n\geq 1$},\\ -\kappa,&\text{if $n=0$},\\ 0&\text{if $n<0$},\end{cases}$ (S41)
@@
$J_{(n)}/J_{0}$ $=$ $\int_{0}^{T}\frac{dt}{T}e^{2\pi int/T}e^{-if(t)+if(t-\delta t)}$ (S42)
$=$ $\frac{1}{\kappa}\int_{C}\frac{dz}{2\pi i}z^{n-1}\frac{(1-\kappa z)(z-\kappa w)}{(z-\kappa)(w/\kappa-z)}$
$=$ $\begin{cases}\kappa^{n}\frac{(1-\kappa^{2})(1-w)}{w-\kappa^{2}},&\text{if $n\geq 1$},\\ 1+\frac{(1-\kappa^{2})(1-w)}{w-\kappa^{2}},&\text{if $n=0$},\\ \left(\frac{\kappa}{w}\right)^{-n}\frac{(1-\kappa^{2})(1-w)}{w-\kappa^{2}}&\text{if $n<0$},\end{cases}$
with $w=e^{2\pi i\delta t/T}$. The coupling to higher Floquet bands mediated by $J_{(n)}$ is always perturbatively small in $\kappa$, so that it can be safely neglected. In turn, the coupling to the first Floquet band for the rescaled case $J_{0}^{\prime}=J_{0}/\kappa$ is of order $1/\kappa$, so that the first Floquet band is accessible to the dynamics and the corrections are on order $TJ_{0}^{2}/\kappa^{2}$. It follows that the first Floquet band can be neglected for $T\ll\kappa^{2}/J_{0}$, in general a more restrictive condition that the one given in Eq. (S38).
SII.2 Working at the sweet spot of flux noise insensitivity.
@@
$a_{l}$ $=$ $4\sqrt{\frac{2}{\pi T\omega}}\int_{-1}^{\kappa}dxx^{l-1}\sqrt{\frac{x}{\kappa-x}+\frac{1}{x\kappa-1}-c},$~~~~~ (S48)
@@
SII.4 Full time-dependent dynamics
In order to check the prediction of the Floquet theory we check the result for the effective model in the fundamental Floquet band and the exact dynamics generated by the time-dependent modulation against the target evolution with constant uniform hopping and a flux $\phi$. In order to keep the analysis simple, we restrict to the single-particle spectrum and solve the dynamics for a set of $N_{s}=4$ transmon. In this subspace, the target Hamiltonian in the frame rotating at frequency $\omega$ reads
@@
where we have gauged the dependence on the external flux between the second and third site [86, 87]. In turn, the evolution in the fundamental Floquet band is described by the effective Hamiltonian
@@
In order to test the goodness of the Floquet modulation, we also study the fidelity $F=|\langle\psi_{\rm target}|P|\psi_{\rm exact}\rangle|$ integrated over the time/phase-shift as a function of time for different values of $T$, while keeping constant ratio $\tau/T=0.3$. The matrix $P={\rm diag}(1,-1,-1,1)$ accounts for the $\pi$ phase shifts due to the Levitons in the external links, as in Eq. (S35). As expected the larger the Floquet frequency $2\pi/T$ the better is the fidelity, as shown in Fig. (S2) where we plot the infidelity 1F1-F for four different values of $T$. For $T=1~{\rm ns}$ we see that the sampling at half period shows oscillations at the Floquet frequency. For higher values of the Floquet frequency we see that the error accumulated grows to few percents after 10 periods of the coherent oscillations set by $J_{0}$.
@@
Table 1: Optimal choice of energy scales (in GHz), partly from Ref. [35].
$\gamma_{\phi}$ $\gamma_{1}$ $\Gamma$ $J_{0}$ $\Omega$ U $1/T$ $1/\tau$ $\omega$
<<0.001 0.001 0.01 0.041 0.04 0.188 2.0 6.67 8.0
SIII Perturbative approach to the many-body spectrum
@@
SIV Master equation numerical approach
@@
Figure S4: Three-photon spectrum for a chain of $N_{s}=4$ sites, as a function of the flux $\Phi$ and the driving frequency $\omega_{d}$, for $J=0.041~{\rm GHz}$, $\Omega=0.04~{\rm GHz}$ and $\Gamma=0.1~{\rm MHz}$, $\omega=3.9~{\rm GHz}$, and $U=0.188~{\rm GHz}$, for different values of the transmon relaxation rate $\gamma_{1}$. (a,b,c) neglecting dephasing, $\gamma_{\phi}=0$; (d,e,f) $\gamma_{\phi}=\gamma_{1}$.
@@
[…]
@@
[…]
Figure S5: Spectrum $|S(\omega_{d})|^{2}$ as a function of the driving frequency $\omega_{d}$ and the flux $\Phi$ for $\Omega=0.04~{\rm GHz}$ in the region frequency around the single-photon and two-photon resonances. The other parameters are as given in the main text and the color bar is made non-linear by setting $z(x,y)^{2/5}$. Note that the range of values covered by the color scale is much smaller than in Fig. 3 of the main text (see also Fig. S5).
@@
To observe flux fractionalization in the $N_{p}=2$ photon sectors we assume low transmon relaxation rate, $\gamma_{1}=1~{\rm MHz}$, low coupling rate to the transmission line resonator, $\Gamma=1\div 10~{\rm MHz}$ and zero transmon pure dephasing rate, $\gamma_{\phi}=0$. As pointed out in the main text, we expect the oscillations in the $N_{p}$ sector soliton to be visible as long as $N_{p}\gamma_{1},\gamma_{\phi}\ll J_{(N_{p})}/N_{s}$. In Fig. S3 we show the fading of the oscillations as we increase the relaxation and dephasing rates. The effective hopping for the $N_{p}=2$ soliton with $J=0.041~{\rm GHz}$ and $U=0.188~{\rm GHz}$ is $J_{(2)}=0.018~{\rm GHz}$, so that in the left panels, where $J_{(2)}/3>2\gamma_{1},\gamma_{\phi}$ oscillations are visible, whereas in the right panels, where $J_{(2)}/3<2\gamma_{1},\gamma_{\phi}$ oscillations are lost.
[…]
@@
In addition, in Fig. S5, we show the plot of the signal for driving $\Omega=0.04~{\rm GHz}$ in the region around the single-photon peak. We see a signal that is strongly modified with respect to the weak driving one. In Fig. S7, top panel, we plot the output signal $S(\omega_{d})$ as a function of the driving frequency $\omega_{d}$ and the driving strength $\Omega$ at zero flux, with specific cuts at different $\Omega$ in the bottom panel. We clearly see that the signal for increased driving strength $\Omega$ is highly modified from the weak driving case. In particular, the single-photon peak splits and transfers its spectral weight to higher photon resonances, that are located at the expected many-body energies. We then see that the two-photon and three-photon peaks are very narrow resonances for weak relaxation rate $\Gamma$, surrounded by larger dips. The peaks are washed out by increasing $\Gamma$, and only dips remain, that nevertheless show the expected evolution with the unperturbed many-body energies, as shown in the main text. In the next section we study the single site driven-dissipative dynamics in presence of Hubbard interaction and ascribe to the combination of driving, dissipation, and interaction the complicate spectrum evolution.
@@
Figure S6: Top: reflection spectrum as a function of the driving frequency $\omega_{d}$ and the driving amplitude $\Omega$, for $\omega=3.9$ GHz, $U=0.188~{\rm GHz}$, $J=0.041~{\rm GHz}$, $\Gamma=0.1~{\rm MHz}$, $\gamma_{1}=0.001~{\rm MHz}$, and we neglect $\gamma_{\phi}$.
SV Single site driven-dissipative dynamics
@@
$\hat{\rho}=\int_{{\cal C}}d\alpha\int_{{\cal C}^{\prime}}d\beta\frac{|\alpha\rangle\langle\beta^{*}|}{\langle\beta^{*}|\alpha\rangle}P(\alpha,\beta)$ (S70)
@@
${\cal I}(a,b,\epsilon)$ $=$ $\int_{{\cal C}}d\alpha\int_{{\cal C}^{\prime}}d\beta\alpha^{c-2}\beta^{d-2}e^{\frac{\epsilon}{\alpha}+\frac{\epsilon^{*}}{\beta}+2\alpha\beta}$ (S73)
@@
We see that in the limit for $U\to 0$ the output field asymptotically approaches the coherent state solution $\langle a^{\dagger}\rangle_{0}=-\Omega/(\omega+i\Gamma/2)$ of a driven-dissipative harmonic oscillator.
In Fig. S7 we plot the output signal as function of the driving strength $\Omega$ and the relaxation rate $\Gamma$. In the top panel we see that as we increase the driving strength the main single-particle peak splits and other peaks at the two-photon and three-photon resonances appear. As the spectral weight is gradually transferred from the single-photon peak to the other resonances a complicate response featuring dips and splittings appears, that is totally due to the combination of driving and interaction. In the bottom panel we show the evolution of the output signal by increasing the relaxation rate. A general intuitive trend appears, in which fine structure details of the signal are washed out by increasing $\Gamma$.
@@
As a check, we have confirmed that the results of Eq. (SV) and the results of the numerical calculation of the steady state for a single site driven-dissipative Bose-Hubbard model coincide.